# SCAR: Scheduling Multi-Model AI Workloads on Heterogeneous Multi-Chiplet Module Accelerators

### v1

@@ IV Scheduling Framework
We discuss our scheduling framework for multi-model workloads on heterogeneous MCMs based on the hierarchical search space characterization and problem formulation in Section III. As illustrated in Figure 2, our scheduling algorithm is a two-level approach: top-level and per-window searches. Top-level search is responsible for selecting layers in each model to be scheduled within a time window and determining the initial number of chiplet nodes for each model. Per-window search explores the spatial and temporal partitioning (i.e. tiles or layer segments) of the layers in each model at the chiplet granularity. To explore the chiplet granularity tiling space, we generate valid inter-chiplet-pipelined schedules utilizing a scheduling tree structure inspired by the RA Tree [4]. Each schedule is evaluated using our custom heterogeneous MCM cost model which provides feedback to the chiplet level tiling (”layer segmentation” in  Figure 2) with expected metrics (latency, energy, EDP, etc.).
We codify our scheduling algorithm into a software framework, as illustrated in Figure 3. As inputs, our scheduling framework receives (1) description files of the multi-model workloads (layer parameters, topology, dependencies, etc.) and (2) a description file of the MCM hardware specification, such as the number of chiplets, the shape, and dataflow organization of the chiplet arrays, NoP bandwidth, on-chiplet memory size, and so on. As outputs, our scheduling framework reports an optimized schedule with expected metrics such as latency, energy, EDP, or other user-defined metrics as a combination of latency and energy. Our scheduling framework consists of four software engines, which handle each step of the scheduler discussed in this section. Each engine is responsible for each step of our two-level scheduling method as illustrated in Figure 2. We discuss each engine and our cost model utilized by the framework next.
@@ IV-A MCM Reconfiguration Engine (MCM-Reconfig)
The MCM-Reconfig engine at the top-level step receives the multi-model workload descriptions with layer information in each model, layer dependency, and expected latency and energy of each layer on each chiplet class offline-analyzed by MAESTRO [24]. The MCM-Reconfig engine is responsible for the window assignment in Figure 2, which (1) generates candidate time window partitioning strategies via sampling a set of discrete points in time reflecting the boundary points between execution windows and (2) assigns layers from models to each time window. As the final assignment of layers to chiplets is not yet known, the decisions in MCM-Reconfig engine are based on expected execution times. Formally, given $|DF|$ dataflow style classes, the expected execution latency for a layer $l$ is:
@@
Time Windows Characterization. MCM-Reconfig engine first specifies the number of windows, which is the coarse-grained scheduling granularity in our scheduling algorithm. We define $n_{splits}$ as a user-defined parameter to characterize the number of time windows and explore proper cut points for each model. For example, in Figure 2, the model A has a cut after layer 6, which led to having layers 1-6 in Window 1. The worst-case latency experienced by a model in the multi-model workload is set as the time horizon which we partition into periodic time windows.
[…]
Greedy Layer Packing Algorithm. Multi-model workload introduces a challenge: the time window boundary determined by the cut points of one model might not be aligned with other models. Therefore, we adopt a first-fit greedy-packing heuristic where layers are assigned to an execution window if their execution time is expected to be within a time window (see Algorithm 1), even if the start and finish time is not aligned with those of a time window. Any layer with execution time lies across two time windows is deferred to the next time window. This approach not only solves the time window - layer execution time misalignment problem, but also facilitates (i) running low-latency layers in earlier windows, preventing starvation of small workloads blocked by heavy workloads. (ii) Dynamically controlling number of time windows when needed by skipping trivial windows without any workloads.
[…]
@@ IV-B Provisioner Engine (PROV)
The PROV engine is responsible for providing an initial estimate on the number of chiplet needed by each model workload in every time window given a candidate partitioning strategy. This assignment is agnostic to the underlying chiplets’ resources or dataflow, and hence we refer to chiplet resources in this state as nodes. We implement our PROV engine for nodes’ distribution across various model workloads using a set of rules. The rules are based on expected latency, energy, EDP, or user-defined metric for each corresponding window. This computational effort is associated with a specified performance optimization goal, denoted as $\mathcal{P}$ where $\mathcal{P}\in\{Lat,Ergy,EDP\}$, Following a uniform distribution rule, the number of nodes $N_{i}$ allocated to the $i^{th}$ model is:
@@
where $\mathbb{E}(\mathcal{P}_{i})$ represents the expected value of the performance optimization goal for model, computed in a manner similar to the expected execution latency formula in Equation (1), whereas the sum term in the denominator represents the sum of all expected values for every model workload in the current time window. Though in principle other allocation strategies can be implemented for the Provisioner, the benefits of having a rule-based Provisioner as such are twofold:
  • The Provisioner becomes specialized in warranting a fair spatial distribution of resources per window across the various model workloads, leaving temporal allocation tasks to be handled through the other engines.
  • Circumventing an additional top-level search space specification with a high degree of uncertainty, acting instead as a regulator mid-way throughout the framework without aggravating the search complexity.
To ensure the progression of all model workloads assigned to a window, we enforce the allocation of at least one resource per model per window to account for rounding errors in Equation (2) when a model workload assigned to the window has negligible computational overhead compared to its peers. The reallocation process iteratively reassigns nodes from models with max number of resources until the constraint is satisfied.
@@ IV-C Segmentation Engine (SEG)
As the first module in the per-window level, the SEG module is instantiated for every time window, receiving topologically sorted sets of layers from each model to be further partitioned into segments. Segmentation is the process of partitioning a set of layers into smaller subsets of layers (i.e., segments or tiles) that can be mapped to a computing resource for exclusive execution throughout the duration of a time window. Different segmentation choices reflect various trade-off points between the layer-sequential and layer-pipelining execution features: the former controls the granularity of layers within each segment to be co-located for sequential execution on a single chiplet resource; the latter exploits inter-layer and -chiplet pipelining opportunities between various segments.
Segmentation Search Space. As per our formulation in Section III, a segmentation candidate is represented by a sequence of splitting points, where candidate splitting points are specified after each layer for each model’s set of layers provided to the SEG Given $|L_{i}|$ and $|N_{i}|$ as the respective number of layers and number of assigned nodes (from PROV) for a model workload $m_{i}$, the max number of segments that can be generated for $m_{i}$ becomes upper bounded by $N_{i}$. Hence, the overall segmentation space complexity is $\mathcal{O}(\Pi_{i}\ |L_{i}|\times{|N_{i}|})$, with the $\Pi_{i}$ indicating the combinatorial space across all models. To aid in managing the rising multi-model segmentation space complexity, we introduce the following heuristics.
Heuristic 1. Product to summation reduction. We enable SEG to navigate the segmentation search space with reduced complexity through a two-step process: (1) SEG leverages the independence of segments from different models to initially explore the segmentation subspace for each model separately (2) Segmentation point candidates from the top-k configurations for each model are used to construct a smaller search space for the combinatorial co-exploration of the segmentation space. Through this heuristic, a sizable reduction in the search space complexity can be realized from $\mathcal{O}(\Pi_{i}\ |L_{i}|\times{|N_{i}|})$ down to $\mathcal{O}(max(|L_{i}|\times{|N_{i}|}))$.
Heuristic 2. Node allocation constraint. We designate an additional user-specified constraint to restrict the number of nodes provisioned to a model within a window based on its number of layers. The constraint is particularly beneficial in cases where there is mismatch between the multi-models’ distribution of computational efforts across their respective layers. For instance, a scenario may arise where a single compute-intensive layer may be assigned to same window alongside dozens of layers from smaller models, leading to an unnecessary explosion in the segmentation search space. The enforcement of this constraint is performed by the PROV.
@@ IV-D Scheduling Engine (SCHED)
The innermost Scheduling engine (SCHED) is responsible for generating the final mapping of layer segments to the physical chiplet accelerators on the target MCM.
Scheduling Search Space. As illustrated in Figure 5, the scheduling search space for the mapping of $M$ model workloads onto $\mathcal{C}$ chiplets can be represented as a forest of scheduling trees. Throughout this sub-section, we use three terms to describe different parts of the scheduling search space: (i) forest; as the entire collection of search trees. (ii) tree; characterizing a single scheduling tree modeling with all the $M$ models involved. (iii) subtree; representing a subset part of each tree exclusively associated with a model $m_{i}\in M$.
Scheduling Tree Composition. every node in a scheduling tree corresponds to a unique chiplet resource on the MCM showcasing its distinctive heterogeneous features (i.e., dataflow). Each chiplet is assigned a unique identifier based on a row-major order traversal across the MCM grid. Tree edges are constructed based on each chiplet’s XY neighbors connected directly through an interposer. Though a node $j$ can be replicated throughout the tree, it can only be visited once, indicating its exclusive occupancy by a model.
Trees Distinction. within each tree, the root nodes of the subtrees specify different chiplets as potential starting positions for candidate model schedules. This is particularly relevant when the underlying pattern of heterogeneity is non-uniform, causing scheduling options to be dependent on the starting position (see Figure 5). Thus, the scheduling space coverage starts by selecting a tree, represented by a permutation sequence of subtrees’ root nodes – e.g., permutation sequence [i,j,k] indicates exploring scheduling candidates for a tree with scheduling candidates starting at chiplet positions $i$, $j$, and $k$ for a 3-model workload. The depth of model $i$’s subtree is determined by its number of provisioned resources $N_{i}$.
Candidate Schedules Generation. Through traversing each subtree, we can obtain candidate execution schedules for each model by assigning segments orderly to the subtree’s nodes. Starting from the root node of the first model’s subtree, a constrained depth first search (DFS) is performed generating a candidate schedule path once the full subtree depth ($N_{i}$) has been reached. This traversal is repeated for each subsequent subtree, constrained on the preceding subtree’s prior visited nodes. Traversal paths from each subtree are then aggregated to form the overall scheduling candidate.
Encoding and Search Algorithm. As shown in Figure 5, use a $2\times|M|$-length tuple to represent the final scheduling encoding, where the first $M$ entries reflect segmentation decisions for each model $m_{i}$, and the latter $|M|$ entries reflect schedule mappings of segments to chiplets for each workload. This form of encoding facilitates supports having different search algorithms for each engine. We tested both brute-force and evolutionary algorithms as will be shown in Section V.
Schedules Starting Positions. We constrain the number of scheduling trees to chiplet nodes that satisfy either of the following two conditions: (i) chiplets that maintain a direct link to an offchip DRAM memory interface (as in the right and left-most chiplets of the MCM in Figure 5). (ii) ending chiplet positions from the preceding window, which enable leveraging data locality across time windows.
@@ IV-E Cost Model and Scoring
TABLE I: MCM microarchitecture parameters from [48, 40]. All numbers are scaled to 28 nm process node technology.
@@
Cost Model. The overall cost model constitutes three distinctive cost model components: offchip communication model, inter-chiplet communication model, and intra-chiplet cost model. We follow our latency and energy modeling characterization in Section III-E and Section III-F, and use the architectural parameters provided in [48, 40] for the offchip and inter-chiplet communication costs as shown in Table I. For the intra-chiplet cost model, we utilize the open-source accelerator cost model, MAESTRO [24, 25], to evaluate intra-chiplet performance based on a chiplet’s underlying dataflow and hardware parameters.
Scoring. Scores are estimated based on latency, energy, or EDP metrics following Section III modeling. The SCHED aggreages scores for each model’s schedule, and returns the top performing configuration to the SEG engine to rank segmentation strategies. Top segmentation strategies in each window are aggregated to score the overall scheduling strategy at MCM-Reconfig (see the scoring flow in Figure 3).
@@ V-A Experimental Settings
TABLE II: Our experimental Multi-model workload scenarios for datacenter and AR/VR use-cases inspired by MLPerf [46, 37] and XRbench [27] benchmarks. ‘sl’ indicates sequence length
@@
Figure 6: The evaluated MCM chiplet organizations.
Multi-Model Workloads. Our evaluations are performed on multi-model workload scenarios based on models from (i) MLPerf inference benchmark [46, 37] for the datacenter multi-tenancy setting; (2) XRBench [27] for AR/VR workloads. The full list of scenarios is provided in Table II covering a wide range of use-cases with varying degrees of diversity and complexity. We implement our MLPerf following datacenter usage trends in [13, 42, 19].
MCM System. We follow Simba’s on-package chiplet arrangement to implement our MCM experimental templates [48]. Simba comprises a total of 36 chiplets connected through a Mesh topology and arranged as four $3\times 3$ groups of chiplets. We implement (1) $3\times 3$, and (2) $6\times 6$ MCM templates for our experiments. For each, we adopt XY routing for on-pacakge data movement, and integrate further memory interfaces on the sides of the outer chiplets, providing direct links to the offchip DRAM via double-sided memory channels as in [10]. We consider 4096 PEs/chiplet and 256 PEs/chiplet for the datacenter and AR/VR settings, respectively. We set the L2 shared memory size in each chiplet to 10 MB, inspired by the on-chip memory size in a recent mobile accelerator [43].
Baselines and Heterogeneity Patterns. We choose Shi-diannao [8] and NVDLA [39] dataflow styles for our accelerator chiplets. We accordingly implement two baselines:
  • Standalone. Each model in a multi-model workload is assigned a single chiplet for execution, all chiplets posses the same dataflow.
  • Simba-like Pipelining. In each time window, Model workloads can be assigned to more than one chiplet to leverage pipelining benefits. All chiplets posses the same dataflow.
We implement several patterns for the heterogeneous on-package integration of Shi-diannao and NVDLA chiplet accelerators. As shown in Figure 6, we test heterogeneous checkerboard, sides, and cross patterns.
@@
Search Algorithms and Evaluation. We adopt a brute-force search for all experiments entailing the $3\times 3$ MCM template. For the scaled $6\times 6$ experiment, we implement an evolutionary algorithm for the SEG module as a meta-heuristic approach to navigate the rising complexity. We set the population size and max number of generations to 10 and 4, respectively. The evaluation criteria follows the scoring function based on hierarchical latency and energy models derived in Section III.
Figure 7: Comparing brute force scheduling space exploration processes across various MCM configurations for various search targets for scenarios 4 and 5 from Table II. standalone Shi-diannao and standalone NVDLA indicate cases where execution is performed on a single chiplet supporting said dataflow.

### v2

@@ IV-A MCM Reconfiguration Engine (MCM-Reconfig)
The MCM-Reconfig engine at the top-level step receives the multi-model workload descriptions with layer information in each model, layer dependency, and expected latency and energy of each layer on each chiplet class offline-analyzed by MAESTRO [35]. The MCM-Reconfig engine is responsible for the window assignment in Figure 3, which (1) generates candidate time window partitioning strategies via sampling a set of discrete points in time as boundary points, and (2) assigns layers from models to each time window. As the final assignment of layers to chiplets is not known apriori, the decisions in MCM-Reconfig engine are based on expected execution times. Formally, given $|DF|$ dataflow style classes, the expected execution latency for a layer $l$ is:
@@
Time Windows Characterization. MCM-Reconfig engine first specifies the number of windows, through a hyperparameter, $n_{splits}$, to explore proper cut points for each model. For example, in Figure 3, the model A has a cut after layer 6, which led to having layers 1-6 in Window 1. The worst-case latency experienced by any model is set as the time horizon to be partitioned into periodic time windows.
Greedy Layer Packing Algorithm. We adopt a first-fit greedy-packing heuristic to assign layers to execution time windows if their execution time is expected to finish within the time window boundaries (see Algorithm 1). Any layer whose execution time lies across two time windows is deferred to the next time window. Through this approach, we enable (i) running low-latency layers in earlier windows (restricts starvation). (ii) dynamically controlling the number of time windows by skipping trivial time windows with no workloads.
Based on our analysis of the periodic window characterization with greedy layer packing using a workload of UNet and GPT2-L against a layer-optimal approach. We found the rate of EDP improvement stagnated after 4 splits. We set $n_{splits}$=4 (5 time windows) as our default unless otherwise stated.
@@ IV-B Provisioner Engine (PROV)
The PROV engine provides an initial estimate on the number of chiplet needed by each model workload in every time window from a candidate partitioning strategy. PROV assignments are agnostic to the underlying chiplets’ properties (dataflow, resources), and hence we refer to chiplets in this state as nodes. We implement the PROV engine to support exhaustive search or rule-based node distribution assignments. A uniform distribution rule allocates $N_{i}$ nodes to the $i^{th}$ model as follows:
@@
where $\mathbb{E}(\mathcal{P}_{i})$ represents the expected value of a target performance optimization metric (latency, energy, EDP) for the model $i$. $\mathbb{E}(\mathcal{P}_{i})$ is computed in a manner similar to the expectation formula in Equation (1).
We ensure every model in the time window is assigned at least one node to progress its execution. The rules enable trading off search complexity for coverage. We analyze the efficacy of the uniform distribution compared to the exhaustive search in Section V.
@@ IV-C Segmentation Engine (SEG)
The SEG module is instantiated every time window to partition topologically sorted model layers into layer segments (Definition 5) that are mappable to computing nodes for exclusive execution throughout the time window. Different segmentation choices reflect various trade-off points between the layer-sqeuencing and layer-pipelining features: the former concerns with execution locality on the same node; the latter specifies inter-layer and -chiplet pipelining opportunities.
Segmentation Search Space. A segmentation candidate is represented by a sequence of splitting points. Candidate splitting points for a model can be specified after each layer provided to the SEG. Given $|L_{i}|$ and $|N_{i}|$ as the respective number of layers and number of assigned nodes from the PROV to model workload $m_{i}$, the max number of segments that can be generated for $m_{i}$ is upper bounded by $N_{i}$. Thus, the overall segmentation space complexity becomes $\mathcal{O}(\Pi_{i}\ \binom{L_{i}}{N_{i}-1})$, We incorporate the following heuristics to manage complexity.
Heuristic 1. Product to summation reduction. We reduce complexity by leveraging the independence of segments from different models to divide the search into a two-step process: (1) SEG first evaluates segmentation candidates from each model separately. (2) The top-k segmentation candidates from each model are used to construct a smaller combinatorial segmentation search space used in the final evlauation. Through this heuristic, the search space complexity is reduced from $\mathcal{O}(\Pi_{i}\ \binom{L_{i}}{N_{i}-1})$ to $\mathcal{O}(max(\binom{L_{i}}{N_{i}-1}))$.
Heuristic 2. Node allocation constraint. To further reduce the search complexity within the SEG engine, SCAR supports having a node allocation constraint as a hyperparameter, particularly beneficial in cases where time windows have model workloads with dissimilar layer distributions (e.g., one model with a heavy layer, another with numerous small layers), which could lead to an explosion in the number of trivial segmentation options from the low-cost layers, causing the the SEG search space complexity to rise. As such small layers are more suited for continuous execution on the same resource by virtue of their lower computational footprints, and to limit unnecessary, costly inter-chiplet data movements, we designate a node allocation constraint to enable alleviating this added complexity by restricting the number of nodes assigned to workloads with disproportionately large number of layers.
@@ IV-D Scheduling Engine (SCHED)
The innermost SCHED engine is responsible for generating the actual physical mapping of layer segments onto chiplets.
SCHED Search Space. As illustrated in Figure 5, the scheduling search space for the mapping of $M$ model workloads onto $\mathcal{C}$ chiplets is represented by a forest of scheduling trees. We define (i) forest; as the entire collection of search trees. (ii) tree; a single scheduling tree with all models. (iii) subtree; a subset tree exclusively associated with a model.
Scheduling Tree Composition. every node in a scheduling tree corresponds to a unique chiplet resource on the MCM showcasing its distinctive heterogeneous features (i.e., dataflow). Tree edges are constructed based on each chiplet’s neighbors connected directly through an interposer. Though a node $j$ can be replicated throughout the tree, it can only be visited once, indicating its exclusive occupancy by a model.
Trees Distinction. within each tree, the root nodes of the subtrees specify different chiplets as potential starting positions for candidate model schedules (see Figure 5). Thus, the scheduling space coverage starts by selecting a tree, represented by a permutation sequence of subtrees’ root nodes – e.g., permutation sequence [i,j,k] indicates exploring scheduling candidates for a tree with scheduling candidates starting at chiplet positions $i$, $j$, and $k$ for a 3-model workload. The depth of model $i$’s subtree is determined by $N_{i}$.
Candidate Schedules Generation. Through traversing each subtree, we can obtain candidate execution schedules for each model by assigning segments orderly to the subtree’s nodes. Starting from the root node of the first model’s subtree, a constrained depth first search (DFS) is performed generating a candidate schedule path once the full subtree depth ($N_{i}$) is reached. This traversal is repeated for each subsequent subtree, constrained on the preceding subtree’s prior visited nodes.
Encoding and Search Algorithm. As shown in Figure 5, use a $2\times|M|$-length tuple to represent the final scheduling encoding, where the first $M$ entries reflect segmentation decisions for each model $m_{i}$, and the latter $|M|$ entries reflect schedule mappings of segments to chiplets for each workload. This encoding enables various search strategies (e.g., evolutionary).
@@ IV-E Cost Model and Scoring
TABLE II: MCM microarchitecture parameters from [64, 55]. All numbers are scaled to 28 nm technology.
@@
Cost Model. We build our analytical cost model on top of MAESTRO [35, 36], which is a standardized analytical cost model tool for modeling performance of AI operators’ mappings (dataflow and tiling) on AI accelerators, with a reported 96% accuracy of performance estimation compared to low-level RTL simulation. MAESTRO natively enables intra-chiplet performance modeling, and we extend it for the MCM AI accelerator setting by integrating the MCM design parameters from Simba [64] as shown in Table II, where we add MCM-specific layers upon MAESTRO leveraging its proven communication overhead modeling methodology to model offchip and inter-chiplet communication costs. The overall cost model follows the performance models in Section III-E.
Scoring. Scores are estimated based on latency, energy, or EDP metrics following the characterization in Section III. The SCHED aggregates scores for each model’s schedule, and returns the top performing configuration to the SEG engine to rank segmentation strategies. Top segmentation strategies in each window are aggregated to score the overall scheduling strategy at MCM-Reconfig (see the scoring flow in Figure 4).
@@ V-A Experimental Settings
TABLE III: Our experimental Multi-model workload scenarios for datacenter and AR/VR use-cases inspired by MLPerf [62, 49] and XRbench [38] benchmarks. ‘sl’ indicates sequence length
@@
Figure 6: Evaluated MCM chiplet organizations in this work.
Multi-Model Workloads. Our evaluations are performed on multi-model workload scenarios based on models from (1) MLPerf inference benchmark [62, 49], which are curated for datacenter multi-tenancy setting following data center usage trends in [23, 58, 29]. (2) XRBench [38] for multi-model AR/VR workloads. The full list of scenarios is provided in Table III covering a wide range of use-cases with varying degrees of diversity and complexity.
MCM System. We evaluate SCAR on a variety of MCM systems following Simba architecture [64]. Simba comprises a total of 36 chiplets arranged as four $3\times 3$ groups of chiplets and connected in a Mesh topology. We implement (1) $3\times 3$, and (2) $6\times 6$ MCM templates for our experiments. we adopt XY routing for on-pacakge data movement, and integrate further memory interfaces on the sides of the outer chiplets for direct access to the offchip DRAM as in [19]. We consider 4096 PEs/chiplet and 256 PEs/chiplet for the datacenter and AR/VR settings, respectively. We set L2 shared memory size in each chiplet to 10 MB as in a recent mobile accelerator’s on-chip memory size [59].
Baselines and MCM patterns. We choose Shidiannao [16] and NVDLA [52] dataflow styles for our accelerator chiplets given their proven superiority [37] and adopt two baselines:
  • Standalone. each model is assigned a single accelerator chiplet - all chiplets adopt the same dataflow.
  • Simba-like Pipelining. In each time window, model workloads can be assigned to more than one chiplet. All chiplets adopt the same dataflow.
Figure 6 illustrates the MCM chiplet patterns evaluated through SCAR. The 3$\times$3 MCM patterns are the default.
@@
Search Criteria. We adopt an exhaustive brute-force search for all $3\times 3$ MCM experiments. For the $6\times 6$ experiment, we implement an evolutionary algorithm for the SEG module to navigate the rising complexity. We set the population size and max number of generations to 10 and 4, respectively.
TABLE IV: Breakdown of the search results across all MLPerf datacenter scenarios from Table III. The comparison involves the top performing models on latency and EDP for each search strategy across all workload scenarios. Latency estinates at 500 MHz.
Strategy Latency Search EDP Search
Latency (s) EDP (J.s) Latency (s) EDP (J.s)
Sc1 Sc2 Sc3 Sc4 Sc5 Sc1 Sc2 Sc3 Sc4 Sc5 Sc1 Sc2 Sc3 Sc4 Sc5 Sc1 Sc2 Sc3 Sc4 Sc5
Stand.(Shi) 1.41 1.41 2.63 14.82 14.82 0.029 0.047 1.21 16.43 21.09 1.41 1.41 2.63 14.82 14.82 0.029 0.047 1.21 16.43 21.0
Stand.(NVD) 0.36 0.36 4.06 7.94 7.94 0.007 0.01 0.786 5.707 7.608 0.36 0.36 4.06 7.94 7.94 0.007 0.01 0.786 5.707 7.608
Simba (Shi) 0.99 0.99 2.44 7.98 8.29 0.02 0.03 1.048 10.85 11.216 1.13 0.97 2.53 8.03 9.79 0.024 0.024 1.097 8.706 13.75
Simba (NVD) 0.25 0.27 1.37 8.34 8.34 0.005 0.007 0.24 6.216 8.225 0.29 0.34 1.36 8.284 8.28 0.006 0.008 0.238 6.165 8.162
Het-CB 0.36 0.39 1.02 6.04 7.39 0.007 0.013 0.331 5.545 9.051 0.51 0.39 1.27 5.6 7.81 0.01 0.013 0.37 5.21 9.166
Het-Sides 0.34 0.34 1.12 3.07 4.81 0.007 .011 0.463 3.074 5.8 0.36 0.38 2.0 3.8 5.19 0.008 0.01 0.35 3.328 6.107
V-B Datacenter multi-model scheduling results
Figure 7: Latency, energy, and EDP evaluations for top-scoring 3$\times$3 candidates normalized by the standalone NVDLA for datacenter scenarios. Highlighted barplots indicate main results (aligned optimization and evaluation metric).
Figure 8: Pareto results for the brute force search across various MCM strategies on various search targets for scenarios 4 and 5 from Table III.
Figure 9: Top-scoring scheduling strategy for Scenario 4 Het-Sides. Allocation of chiplets and coarse-grained schedules are shown for each time window. Times computed over 500 MHz indicate cumulative window latencies.
We discuss the full datacenter scheduling results on the $3\times 3$ MCM, provided in Table IV and Figure 7. Figure 8 illustrates the Pareto search results for Scenarios 3 and 4.
Impact of Workload Scenarios. We observe that different scenarios are more affine towards different MCM strategies owing to their varying degrees of diversity and computational demands within the multi-model suite. For example in the EDP search in Table IV, scenarios 1-3 favor the standalone (NVD) and Simba (NVD) strategies as both are dominated by low-batched, transformer-based workloads (GPT-L and BERT-L) characterized by strong affinities towards the NVDLA dataflow style, As the computational load and model diversity increase, schedules obtained through heterogeneous strategies become more favorable. For instance in the EDP search scenarios 4-5 with more batches and heavy ResNet-50 and UNet workloads, Het-Sides outperforms all other strategies on EDP evaluation, experiencing 46.02% and 25.18% less EDP compared to Simba (NVD) on the respective scenarios.
Target Objective Impact. In matching criteria plots (A1, B2, and C3) from Figure 7, Het-Sides schedules for the heaviest scenarios 4-5 outperform schedules from all other strategies. This superiority, however, comes at the expense of other metrics. For instance in C3, the Het-Sides schedule outperforms that from all other strategies by experiencing 0.58$\times$ the EDP from the Standalone (NVD). This is achieved, however, by speeding up execution at the expense of energy consumption, where Het-Sides achieves the fastest execution speedup of 2$\times$ over Standalone (NVD) in (A3) at the expense of being 1.22×1.$2$ more energy demanding than Standalone (NVD) in (B3).
Pipelining Benefits and NoP impact. Pipelining individually can offer performance improvements over standalone baselines when ample resources are available to the models. For example, Scenario 3 (Latency Search) – top-left most Pareto in Figure 8 – shows Simba (NVD) schedules achieving considerable speedups over the the standalone NVDLA baseline (2.9 $\times$ for the top-performing candidate from Table IV). The reason being that as the scheduler targets latency optimization, it attempts to consistently identify chiplet assignment strategies that equate pipelining latencies between model workloads assigned to the same time window. For instance, in the last time window containing ResNet-50, the Simba (NVD) scheduler for Scenario 3 queued 18, 14, and 29 layers from the GPT-L, BERT-L, and ResNet-50 for processing, provisioned 3 chiplets for each subset, and yielded accordingly 3 layer segments per model. This configuration led to the model workloads achieving comparable inter-chiplet pipelining latencies of 0.28 ms, 0.32, and 0.33 ms, eventually leading the full schedule to sustain a total of 1.37 s latency compared to the 4.06 s from the standalone NVDLA. We also observe that keeping medium-sized data traffic on package through chiplet-to-chiplet data passing contributes additional energy savings for these medium sized workloads by decreasing offchip memory read/write accesses at the segments’ start and end times.
For the NoP traffic, Figure 7-A1 summarizes our findings. In Scenarios 1-3, Simba (NVD) leverages inter-chiplet pipelining to outweigh added NoP costs, achieving latency speedups over standalone NVDLA reaching 1.4×1.4$\times$1.4 ×, 1.3×1.3$\times$1.3 ×, and 2.9×2.9$\times$2.9 ×, respectively. In Scenarios 4-5, Simba (NVD) to incur 1.05×1.05$\times$1.05 × slowdown from the added traffic in both scenarios.
Heterogeneity Benefits. Heterogeneous integration patterns compensate for the added NoP traffic in the heavier workloads (Scenarios 4-5) and boost efficiency through considering per-layer dataflow affinities. Het-Sides achieves 1.7$\times$ and 1.25×1.25$\times$1.25 × EDP efficiency over the standalone (NVD). We find also that Het-Sides is always superior to Het-CB in such heavy scenarios as it offers opportunities for both homogeneous and heterogeneous inter-chiplet pipelining. This is beneficial for batched layer sequences with same dataflow affinities.
Het-Sides Top Schedule. In Figure 9, we illustrate the top-scoring Het-Sides schedule from the EDP search in Scenario 4. As shown, the greedy-packing algorithm leads to non-uniform windows where smaller workloads (ResNet-50) are assigned to the earlier windows at the expense of larger workloads (BERT-L). This facilitates (i) fine grained optimization of small workload schedules; (ii) avoiding small workloads starvation. GPT-L and BERT workloads dominate the schedule from window 2. Table VI breaks down the latency in each window.
V-C AR/VR multi-model scheduling results
The AR/VR scheduling results on the EDP search for the $3\times 3$ MCM are provided in Table V and Figure 10. Figure 11 illustrates their the Pareto search results. On average, the Het-Sides schedule achieves $17\%$ and $21.3\%$ EDP improvement over standalone NVDLA and Simba (NVD), respectively. We also find that the smaller MCM system limits improvement potential for the heaviest scenarios, AR assistant (scenario 6) and AR gaming (scenario 7) due to increased resource contention, leading the heterogeneous and NVDLA based schedules to achieve comparable evaluations (see Figure 10).
Figure 10: Evaluations on the EDP search for the XRBench usage scenarios listed in Table 1 normalized by NVDLA standalone configuration.